Please help I'm stuck on a linear equation using the ELIMINATION method.
The question is as follows:
2x+3y =7
3x-2y =4
Would appreciate any help.

Answer by rothauserc(4718)

2x+3y =7
3x-2y =4
multiply the first equation by 2 and multiply the second equation by 3
4x+6y =14
9x-6y =12
add the resulting equations
13x =26
x = 2
substitute for x in first equation
2*2 +3y = 7
3y = 3
y = 1
check the answers by substituting the second equation
3*2 -2*1 =4
6 -2 = 4
4 =4
answers check
x = 2, y = 1